OpenStudy (anonymous):

solve: (x-5)^2+(x-5)-6=0

OpenStudy (imstuck):

You will have to expand the (x-5)^2 and get this:

OpenStudy (imstuck):

\[x ^{2}-10x +25+x-5-6=0\]then combine like terms.

OpenStudy (imstuck):

\[x ^{2}-9x+14=0\]Then factor that. Yes?

OpenStudy (imstuck):

That factors very nicely to (x-2)(x-7)=0

OpenStudy (imstuck):

And there you have it!

OpenStudy (anonymous):

or you can just solve it without multplying out \[ (x-5)^2+(x-5)-6=0\] \[(x-5-2)(x-5+3)=0\] etc